Evaluate the integral "\oint" 1/(z2+1)(z2-4)dz where |z|=1.5


1
Expert's answer
2021-09-13T17:09:54-0400

First, we suppose that the circle "|z|=1.5" is oriented counterclockwise, as the orientation was not specified. Second, as the function "f(z)=\frac{1}{(z^2+1)(z^2-4)}" is meromorphic inside the disc "|z|\leq 1.5", we can represent this integral as the sum of residues inside the disc :

"\int_\mathcal{C} \frac{1}{(z^2+1)(z^2-4)}dz=2\pi i\sum_{|a|<1.5}Res_a \frac{1}{(z^2+1)(z^2-4)}dz"

The residues of a meromorphic function are its poles, and the poles inside the disc are "z_1=i, z_2=-i" (the poles are the zeros of the denominator with "|z|<1.5" ). As both poles are simple, we can calculate the respective residues using the formula "Res_af(z)dz=\lim_{z\to a} (z-a)f(z)" :

"Res_i f(z)dz=[\frac{1}{(z+i)(z^2-4)}]_{z=i}=\frac{i}{10}"

"Res_{-i}f(z)dz=[\frac{1}{(z-i)(z^2-4)}]_{z=-i}=\frac{-i}{10}"

Therefore, the intergal is "\int_\mathcal{C} \frac{1}{(z^2+1)(z^2-4)}dz=0" and is, in fact, independent of the orientation.